OpenStudy (anonymous):

6 over 35 r = 18 over 25

OpenStudy (baldymcgee6):

6/35 * r = 18/25 6r = 126/5 r= 21/5

Directrix (directrix):

6/(35r) = 18/25 r = 5/21 ---->* ------------------ Note to @jazy0512 --> It is important to post questions so that it is clear which terms are in the numerator and which are in the denominator.